# algorithmic

### v4

@@ 4.1 Almost Equivariant Models
Definition 4.1 ($\varepsilon$-almost equivariant model).
@@
[…]
Figure 2: We provide a visualization of the grid used to perform Lie algebra convolutions. In this example, we sample $n=10$ elements, $x_{0},\ldots,x_{9}$, of the Lie algebra $\mathfrak{so}(2)$ and apply $\Phi$, parameterized as a single-layer feedforward neural network, $\mathcal{N}$, to acquire the grid elements, $g_{0}=\mathcal{N}^{-1}(x_{0}),\ldots,g_{9}=\mathcal{N}^{-1}(x_{9})$, that are convolved over. The $\times$ symbol shows the action of the rotation on the grid’s lower-right corner. The linearity of both the Lie algebra and the feedforward neural network are apparent here, although the neural network’s activation function introduces subtle non-linear effects on the grid structure. We experimented with normalizing the grid values to the range $[-1,1]$ along both the $x$ and $y$ axes but found it did not lead to reliable improvements in performance. We suspect that the variance in scale across grids might help the model learn functions across a range of scales.
@@
4.2 Lie Algebra Convolutions
Similar to the approach taken in van der Ouderaa et al. (2022), we build an almost equivariant neural network layer based on the Lie algebra, $\mathfrak{g}$, of a matrix Lie group, $G\subseteq GL_{n}(\mathbb{R})$. However, our model makes use of a few, key differences. First, rather than parametrizing our kernel in a finite-dimensional random Fourier features basis, we instead encode the Lie algebra basis explicitly. For most matrix Lie groups, the corresponding Lie algebra basis has an easily calculated set of generators, i.e. a set of basis elements $\{x_{i}\}$. Second, instead of mapping elements of $\mathfrak{g}$ directly to $G$ via the exponential map, we train a neural network, $\mathcal{N}_{\theta}:\mathfrak{g}\to\mathbb{R}^{n\times n}$, to learn an approximation to this mapping directly from data. In our experiments, each Lie algebra convolutional layer of the network is equipped with its own $\mathcal{N}$, which is parameterized as an MLP with a single linear layer followed by a non-linear activation, either a ReLU or a Sigmoid function. Our method confers some key benefits over previous approaches. For one, the kernels used in some past works are still constrained to take as input only group elements, $u,x\in G$, which to some extent limits the flexibility with which they can model partial equivariances. In contrast, our kernel can take any $u,x\in\mathbb{R}^{n\times n}$ as an input, allowing us to model a more flexible class of functions while still maintaining the interpretability achieved by parameterizing this function class via elements of the Lie algebra.
Furthermore, whereas van der Ouderaa et al. (2022) relax equivariance constraints by letting their kernel depend on an absolute group element, $v$, we define a simpler convolution that still allows us to relax equivariance constraints.
Definition 4.2 (Almost Equivariant Lie Algebra Convolution).
[…]
Here, instead of integrating with respect to the Haar measure, we instead integrate with respect to the Lebesgue measure, $\mu$, defined on $\mathbb{R}^{n\times n}$. This is possible because we are integrating over the Lie algebra, $\mathfrak{g}$, which is a vector subspace of $\mathbb{R}^{n\times n}$. Existing works require integrating with respect to the Haar measure because it is finite for compact groups, which allows one to more easily do MCMC sampling. Compactness is also necessary to define fully-equivariant group convolutions parameterized in the Lie algebra, because such a parameterization relies on the exponential map being surjective. Furthermore, while MacDonald et al. (2022) define a method for sampling from the Lie group, $G$, that allows the group convolution to retain full equivariance, even for non-compact groups, by using a measure induced on the Lie algebra by the Haar measure, we adopt our simpler approach since we are not aiming for full group equivariance and instead only for almost equivariance. Thus, we use a uniform measure on the Lie algebra, which for the groups studied here amounts to the Lebesgue measure on $\mathbb{R}^{n\times n}$. While we still ultimately convolve with group elements (in the case of compact groups, for which $\exp:\mathfrak{g}\to G$ is surjective), our inputs, $u$, are taken from the Lie algebra, $\mathfrak{g}$, and then pushed onto the Lie group, $G$, via the $\exp$ map.
Additionally, because the $\exp$ map is surjective only for compact Lie groups (Hall, 2015), the approach of parameterizing Lie group elements by applying the $\exp$ map to elements of the Lie algebra only works in the compact case. Because we model the mapping function $\mathcal{N}_{\theta}:\mathfrak{g}\to G$ using a neural network, our approach extends to non-compact Lie groups.
Finally, our approach easily interpolates between full equivariance, partial equivariance, and non-equivariance. When presented with fully equivariant training data, our neural network over Lie algebra elements can learn the exponential map. When presented with almost equivariant training data, this same neural network can learn an approximation to the exponential map that is justified by said data. And finally, when presented with a task for which equivariance is not beneficial, the neural network is free to learn an arbitrary function over the Lie algebra that best models the training data.
@@ 4.2 Lie Algebra Convolutions
Table 1: Rot-MNIST classification accuracies and RMSE prediction errors for pendulum trajectory prediction. The first column gives the Lie group with respect to which equivariance or almost equivariance (depending on the model) is imposed. The second column gives the number of Lie group/algebra elements sampled by the model, if applicable. Best results are bold-faced and second-best are colored gray.
4.3 Computational Considerations
Calculating convolutions over continuous groups is a non-trivial computational problem as it involves taking an integral which can only be numerically approximated on a finite-precision computer. One approach is to discretize the underlying group and compute a finite sum, but this actually leaves the model layer equivariant only to the action of the group elements in the discretization. MacDonald et al. (2022) introduced a Markov Chain Monte Carlo (MCMC) method which uses the Haar measure to sample from probability distributions over an arbitrary Lie group, thereby enabling fast and easily parallelizable approximations of group convolutions. A similar approach is taken in Finzi et al. (2020), although their method is restricted to Lie groups for which the exponential map is surjective.
We proceed in a similar manner, by defining a uniform probability distribution with respect to the Lebesgue measure over the Lie algebra, $\mathfrak{g}$, and drawing samples to create a grid of Lie algebra elements. These samples are then passed through $\mathcal{N}$, inverted, and convolved over. We provide a visualization of this grid in Figure 2.
@@ 5 Results
We test our Almost Equivariant G-CNN on a suite of tasks that span the gamut of full and almost equivariance. For each task, we compare the performance of our model with that of the Residual Pathway Prior model given in Finzi et al. (2021b), the Approximately Equivariant G-CNN defined in Wang et al. (2022a), the $E(2)$-equivariant and steerable E2CNN of Weiler & Cesa (2019), and a Standard CNN that is equivariant only to translations of the inputs.
@@ 5.1 Image Classification
We first test our model on an image classification task. We focus on the Rot-MNIST dataset, which consists of images taken from the MNIST dataset and subjected to random rotations. We would expect rotational equivariance to be beneficial for classifying these images. The training, validation, and test sets contain 10,000, 2,000, and 50,000 images, respectively. We summarize our results on this task in Table 1. We perform a comprehensive hyperparameter grid search during training, and find that our best-performing model significantly outperforms the almost equivariant baselines that we tested against. It also outperforms the standard CNN and is marginally outperformed only by the fully-$E(2)$-equivariant E2CNN. We didn’t perform any optimization of the kernel functions for any of the models, nor of the neural network mapping from the Lie Algebra to the Lie Group for our model, and expect that with further hyperparameter tuning as well as deeper models and more complex kernel functions, we could achieve even higher performance(s) on the test set. We provide further details on the model training process in the Appendix.
@@ 5.3 Smoke Plume
Next, we test our model on an almost equivariant prediction task. The dataset we use is the smoke plume dataset of Wang et al. (2022a) consisting of $64\times 64$ 2D velocity vector fields of smoke simulations with different initial conditions and external forces, all generated using the PDE simulation framework, PhiFlow (Holl et al., 2020). Specifically, we use the subset of the data that features rotational almost equivariance. As stated in Wang et al. (2022a), “both the inflow location and the direction of the buoyant forces possess a perfect rotation symmetry with respect to the $C_{4}$ group, but the buoyancy factor varies with the inflow positions to break the rotational symmetry.” All models are trained to predict the raw velocity fields at the next time step given the raw velocity fields at the previous timestep as input.
Due to computational constraints, we only run our method on this data and compare to the baseline results reported in Wang et al. (2022a). Table 2 shows how our method compares to the baselines. Due to computational constraints, we were unable to run a full hyperparameter sweep and suspect that doing so would boost our model’s performance even further.
@@ 5.4 Jet Flow
Finally, we test on one more almost equivariant dataset. As described in Wang et al. (2022a), this dataset contains samples of 2D turbulent velocity fields taken from NASA multi-stream jets that were measured using time-resolved particle image velocimetry as described in Bridges & Wernet (2017). We follow the procedure described in Wang et al. (2022a), and “train and test on twenty-four $62\times 23$ sub-regions of jet flows.” Table 2 shows our results.
Group Num Samples Model Jet Flow (RMSE) Smoke Plume (RMSE)
Future Domain Future Domain
E(2) 10 E2CNN $0.1919\pm 0.0016$ $0.2131\pm 0.0023$ 1.05 0.76
10 Residual Pathway Prior $0.1947\pm 0.0066$ $0.2143\pm 0.0057$ 0.96 0.83
T(2) N/A Standard CNN $0.2109\pm 0.0068$ $0.2218\pm 0.0008$ 1.21 1.10
Table 2: Prediction RMSE on simulated smoke plume velocity fields and jet flow 2D turbulent velocity fields with almost rotational symmetry. The results for the baseline methods are taken from Wang et al. (2022a) and compared against our Almost Equivariant G-CNN. As stated in Wang et al. (2022a), Future prediction involves testing on data that lies in the future of the training data. Domain prediction involves training and test data that are from different spatial domains. Best results are bold-faced and second-best are colored gray.
@@ 6 Discussion
In this work, we proposed a definition of almost equivariance that encompassed previous definitions of full and approximate/partial/soft equivariance. We connected this definition to mathematical theory by showing that, given an abelian isometry group, $G$, acting on a Riemannian manifold, $M$, then any isometry, $f$ of $M$, is equivariant to the action of $G$, and furthermore that there exists an $\varepsilon$-almost isometry, $f_{\varepsilon}$ of $M$, not more than $\varepsilon$ from $f$ in the supremum norm, such that $f_{\varepsilon}$ is almost equivariant to the action of $G$. Next, we showed that nothing is lost by taking $f$ and $f_{\varepsilon}$ to be isometric and almost isometric embeddings, respectively, of $M$ into $\mathbb{R}^{n}$. We then appealed to Ulam Stability Theory and Fickett’s Theorem to give conditions under which we can get an isometry of a complete, real Hilbert space close to an almost isometry of the same space. All of this taken together demonstrates that there exist almost equivariant functions that are never “far” from fully equivariant functions, given some constraints on the group action and class of functions, in a sense that can be mathematically quantified.
We next introduced a convolution on the elements of a Lie algebra, for which Lie algebra elements are sampled using the Lebesgue measure on the algebra, that approximates a fully equivariant group convolution. We then showed that such a convolution can model almost equivariance relative to any group action, even those of non-compact groups. We validated our assumptions by testing our model on a 2D image classification task, a 1D sequence regression task, and a 2D sequence regression task. On all tasks, our model exceeded or met the performance of state-of-the-art equivariant and almost equivariant baseline models. This demonstrates the utility of our method across a variety of scientific domains and prediction task types.
@@ 7 Future Work
One line of future work will involve testing our model architecture on a wider class of group actions. While our model is general enough to handle the action of any group, including those of non-compact groups, we have not yet tested it on groups aside from $E(2)$. One reason for this is that the utility of imposing equivariance to non-compact groups or non-Euclidean isometry groups is less clear than for compact Lie groups that are isometries of $\mathbb{R}^{n}$, although Lawrence & Harris (2023) points to some potential applications. We plan to address the case of non-compact groups in a subsequent paper.
Next, there exist a number of ways to further expound upon the theoretical results given here. One potential angle to consider is whether variations of Theorems 3.11 and 3.12 can be made to hold for arbitrary Riemannian manifolds and not just Hilbert and Euclidean spaces, respectively. Another direction would involve undertaking a rigorous analysis of the conditions under which almost equivariance to the action of a non-abelian group can be imposed upon a function. We here gave proof of the existence of almost isometries of Riemannian manifolds that are almost equivariant to certain abelian group actions, which we believe to be the most useful direction as, in practice, one normally seeks to take a fully equivariant model and make it almost equivariant. That said, the more difficult mathematical question is to consider when, given an almost equivariant function on a manifold, it can be transformed into a fully equivariant function on the same manifold. We leave this direction for future work.
Finally, it is known that fully-equivariant kernel sharing for G-CNNs requires that the group act transitively on the input space (Weiler et al., 2021). An interesting direction for future work would be investigating the extent to which this assumption is required for almost equivariant kernel sharing.
@@ A.1.1 Almost Equivariance of the Lie Algebra Convolution
See 3.11 This proof, taken from (Hyers & Ulam, 1945), is reproduced here for the reader’s convenience.
@@ A.2 Mathematical Background
We give brief introductions to the subjects of representation theory, differential topology and geometry, and Lie theory, stating only those definitions and theorems needed to understand the paper. For more comprehensive background, we encourage readers to consult any of Fulton & Harris (2004); Etingof et al. (2011); Hall (2015) for representation theory, any of Lee (2003; 2018) for differential topology and geometry, and Hall (2015) for Lie theory.

### v6

@@ 4.1 Almost Equivariant Models
Definition 4.1.
@@
Note that our definition naturally encompasses full equivariance with respect to the action of connected, compact Lie groups, for which the $\exp$ map is surjective, and which occurs when we take $\varepsilon=0$ and define $\Phi:=\exp$.
@@
Table 1: Rot-MNIST classification accuracies and RMSE prediction errors for pendulum trajectory prediction. Best results are bold-faced and second-best are colored gray.
4.2 Lie Algebra Convolutions
We build an almost equivariant neural network layer based on the Lie algebra, $\mathfrak{g}$, of a matrix Lie group, $G\subseteq GL_{n}(\mathbb{R})$. To parameterize our kernel function, we encode the Lie algebra basis explicitly. For most matrix Lie groups, the corresponding Lie algebra basis has an easily calculated set of generators, i.e. a set of basis elements $\{x_{i}\}$. Second, instead of mapping elements of $\mathfrak{g}$ directly to $G$ via the exponential map, we train a neural network, $\mathcal{N}_{\theta}:\mathfrak{g}\to\mathbb{R}^{n\times n}$, to learn an approximation to this mapping directly from data. In our experiments, each Lie algebra convolutional layer of the network is equipped with its own $\mathcal{N}$, which is parameterized as an MLP with a single linear layer followed by a non-linear activation, either a ReLU or a sigmoid function. Our method confers some key benefits over previous approaches. For one, the kernels used in some past works are still constrained to take as input only group elements, $u,x\in G$, which to some extent limits the flexibility with which they can model partial equivariances. In contrast, our kernel can take any $u,x\in\mathbb{R}^{n\times n}$ as an input, allowing us to model a more flexible class of functions while still maintaining the interpretability achieved by parameterizing this function class via elements of the Lie algebra.
Definition 4.2.
[…]
Here, instead of integrating with respect to the Haar measure, we instead integrate with respect to the Lebesgue measure, $\mu$, defined on $\mathbb{R}^{n\times n}$. This is possible because we are integrating over the Lie algebra, $\mathfrak{g}$, which is a vector subspace of $\mathbb{R}^{n\times n}$. Existing works require integrating with respect to the Haar measure because it is finite for compact groups, which allows one to more easily do MCMC sampling. Compactness is also necessary to define fully-equivariant group convolutions parameterized in the Lie algebra, because such a parameterization relies on the exponential map being surjective. Furthermore, while MacDonald et al. [31] define a method for sampling from the Lie group, $G$, that allows the group convolution to retain full equivariance, even for non-compact groups, by using a measure induced on the Lie algebra by the Haar measure, we adopt our simpler approach since we are not aiming for full group equivariance and instead only for almost equivariance. Thus, we use a uniform measure on the Lie algebra, which for the groups studied here amounts to the Lebesgue measure on $\mathbb{R}^{n\times n}$. While we still ultimately convolve with group elements (in the case of compact groups, for which $\exp:\mathfrak{g}\to G$ is surjective), our inputs, $u$, are taken from the Lie algebra, $\mathfrak{g}$, and then pushed onto the Lie group, $G$, via the $\exp$ map.
Additionally, because the $\exp$ map is surjective only for compact Lie groups [18], the approach of parameterizing Lie group elements by applying the $\exp$ map to elements of the Lie algebra only works in the compact case. Because we model the mapping function $\mathcal{N}_{\theta}:\mathfrak{g}\to G$ using a neural network, our approach extends to non-compact Lie groups.
Group Num Samples Model Jet Flow (RMSE) Smoke Plume (RMSE)
Future Domain Future Domain
E(2) 10 E2CNN $0.1919\pm 0.0016$ $0.2131\pm 0.0023$ 1.05 0.76
10 Residual Pathway Prior $0.1947\pm 0.0066$ $0.2143\pm 0.0057$ 0.96 0.83
T(2) N/A Standard CNN $0.2109\pm 0.0068$ $0.2218\pm 0.0008$ 1.21 1.10
Table 2: Prediction RMSE on simulated smoke plume velocity fields and jet flow 2D turbulent velocity fields with almost rotational symmetry. The results for the baseline methods are taken from Wang et al. [44] and compared against our Almost Equivariant G-CNN. As stated in Wang et al. [44], Future prediction involves testing on data that lies in the future of the training data. Domain prediction involves training and test data that are from different spatial domains. Best results are bold-faced and second-best are colored gray.
Theorem 4.3.
[…]
is $\varepsilon$-almost equivariant.
We provide the proof in the appendix.
@@ 5 Results
We test our Almost Equivariant G-CNN on a suite of tasks that span the gamut of full and almost equivariance. For each task, we compare the performance of our model with that of the Residual Pathway Prior model given in Finzi et al. [12], the Approximately Equivariant G-CNN defined in Wang et al. [44], the $E(2)$-equivariant and steerable E2CNN of Weiler and Cesa [46], and a Standard CNN that is equivariant only to translations of the inputs.
@@ 5.1 Image Classification
We first test our model on an image classification task. We focus on the Rot-MNIST dataset, which consists of images taken from the MNIST dataset and subjected to random rotations. We would expect rotational equivariance to be beneficial for classifying these images. The training, validation, and test sets contain 10,000, 2,000, and 50,000 images, respectively. We summarize our results on this task in Table 1. We perform a comprehensive hyperparameter grid search during training, and find that our best-performing model outperforms all baselines that we tested against. It also outperforms the standard CNN and is marginally outperformed only by the fully-$E(2)$-equivariant E2CNN. We didn’t perform any optimization of the kernel functions for any of the models, nor of the neural network mapping from the Lie Algebra to the Lie Group for our model, and expect that with further hyperparameter tuning as well as deeper models and more complex kernel functions, we could achieve even higher performance(s) on the test set. We provide further details on the model training process in the appendix.
@@ 5.3 Smoke Plume
Next, we test our model on an almost equivariant prediction task. The dataset we use is the smoke plume dataset of Wang et al. [44] consisting of $64\times 64$ 2D velocity vector fields of smoke simulations with different initial conditions and external forces, all generated using the PDE simulation framework, PhiFlow [19]. Specifically, we use the subset of the data that features rotational almost equivariance. As stated in Wang et al. [44], “both the inflow location and the direction of the buoyant forces possess a perfect rotation symmetry with respect to the $C_{4}$ group, but the buoyancy factor varies with the inflow positions to break the rotational symmetry.” All models are trained to predict the raw velocity fields at the next time step given the raw velocity fields at the previous timestep as input.
Due to computational constraints, we only run our method on this data and compare to the baseline results reported in Wang et al. [44]. Table 2 shows how our method compares to the baselines. Due to computational constraints, we were unable to run a full hyperparameter sweep and suspect that doing so would boost our model’s performance even further.
@@ 5.4 Jet Flow
Finally, we test on one more almost equivariant dataset. As described in Wang et al. [44], this dataset contains samples of 2D turbulent velocity fields taken from NASA multi-stream jets that were measured using time-resolved particle image velocimetry as described in Bridges and Wernet [2]. We follow the procedure described in Wang et al. [44], and “train and test on twenty-four $62\times 23$ sub-regions of jet flows.” Table 2 shows our results.
@@ 6 Discussion
In this work, we proposed a definition of almost equivariance that encompassed previous definitions of full and approximate/partial/soft equivariance. We connected this definition to mathematical theory by showing that, given an abelian isometry group, $G$, acting on a Riemannian manifold, $M$, then any isometry, $f$ of $M$, is equivariant to the action of $G$, and furthermore that there exists an $\varepsilon$-almost isometry, $f_{\varepsilon}$ of $M$, not more than $\varepsilon$ from $f$ in the supremum norm, such that $f_{\varepsilon}$ is almost equivariant to the action of $G$. Next, we showed that nothing is lost by taking $f$ and $f_{\varepsilon}$ to be isometric and almost isometric embeddings, respectively, of $M$ into $\mathbb{R}^{n}$. We then appealed to Ulam Stability Theory to give conditions under which we can get an isometry of a complete, real Hilbert space close to an almost isometry of the same space. All of this taken together demonstrates that there exist almost equivariant functions that are never “far” from fully equivariant functions, given some constraints on the group action and class of functions, in a sense that can be mathematically quantified.
We next introduced a convolution on the elements of a Lie algebra that approximates a fully equivariant group convolution. We then showed that such a convolution can model almost equivariance relative to any group action, even those of non-compact groups. We validated our assumptions by testing our model on a 2D image classification task, a 1D sequence regression task, and a 2D sequence regression task. On all tasks, our model exceeded or met the performance of state-of-the-art equivariant and almost equivariant baseline models. This demonstrates the utility of our method across a variety of scientific domains and prediction task types.
@@ 7 Future Work
One line of future work will involve testing our model architecture on a wider class of group actions. While our model is general enough to handle the action of any group, including those of non-compact groups, we have not yet tested it on groups aside from $E(2)$. Lawrence and Harris [27] points to some potential applications of equivariance to non-compact Lie groups.
Next, there exist a number of ways to further expound upon the theoretical results given here. One potential angle to consider is whether variations of Theorem 3.11 can be made to hold for arbitrary Riemannian manifolds and not just Hilbert and Euclidean spaces, respectively. Another direction would involve undertaking a rigorous analysis of the conditions under which almost equivariance to the action of a non-abelian group can be imposed upon a function. We here gave proof of the existence of almost isometries of Riemannian manifolds that are almost equivariant to certain abelian group actions, which we believe to be the most useful direction as, in practice, one normally seeks to take a fully equivariant model and make it almost equivariant. That said, the more difficult mathematical question is to consider when, given an almost equivariant function on a manifold, it can be transformed into a fully equivariant function on the same manifold. We leave this direction for future work.
Finally, it is known that fully-equivariant kernel sharing for G-CNNs requires that the group act transitively on the input space [47]. An interesting direction for future work would be investigating the extent to which this assumption is required for almost equivariant kernel sharing.
8 Acknowledgements
We thank Frederic Sala, Jason Hartford, and Andrew Zimmer for their valuable feedback on this work.
@@ A.1.1 Almost Equivariance of the Lie Algebra Convolution
See 3.11 This proof, taken from [22], is reproduced here for the reader’s convenience.
@@ A.2 Mathematical Background
We give brief introductions to the subjects of representation theory, differential topology and geometry, and Lie theory, stating only those definitions and theorems needed to understand the paper. For more comprehensive background, we encourage readers to consult any of Fulton and Harris [13], Etingof et al. [8], Hall [18] for representation theory, any of Lee [28, 29] for differential topology and geometry, and Hall [18] for Lie theory.
@@ A.3.2 Rotated MNIST Classification
A.4 Computational Resources
All experiments were conducted on a single NVIDIA A100 GPU with 80GB of memory.